Fermi National Laboratory
 

Accelerator Update for November 10, 2006 - November 13, 2006

 

Friday November 10

The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring a Recycler stash (117.9E10) and the Antiproton Source (Pbar) stack (56.0E10), with Linac, Booster, and the Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for a store, and with MiniBooNE, and NuMI taking beam.

At 9:33 AM, Operations was about to load antiprotons into the TeV for the next store, but the Recycler had an emittance blow up. The TeV conducted studies while Recycler experts investigated the problem.

At 9:54 AM, Operations reported that a TeV RF station (TRF8) had no output. Operations contacted experts.

Recycler experts reported at 11:09 AM that they had repaired a damper problem.

Operations prepared CDF for an access at 11:49 AM.

CDF came out of access and Operations put the TeV through a dry squeeze at 1:30 PM.

Operations put the TeV into shot setup at 2:02 PM.

Operations established store 5062 at 4:17 PM with an initial luminosity of 128.19E30.

The Recycler completed its fourth and last stash for the day at 10:50 PM.

Saturday November 11

The midnight shift began with Operations monitoring store 5062 (51.81E10) and a stash (113.2E10), with Pbar stacking (37.1E10), and with NuMI and MiniBooNE taking beam.

At 9:23 AM, cryo system technicians reported that the TeV sector A3 wet engine needed some valve adjustments.

At 11:24 AM, Operations reported that the Booster’s chopper (B:CHOP) had been failing. Operations called in experts.

At 12:34 PM, the Booster experts determined that the B:CHOP problem was in the tunnel. Operations prepared Linac for an access.

At 1:41 PM, Operations reported a problem with a Linac RF station (KRF3). Operations called in an expert.

Booster experts reported at 1:42 PM that they connected a new cable for B:CHOP.

Operations terminated store 5062 at 4:15 PM.

Operations put the TeV into shot setup at 4:25 PM.

Operations established store 5063 at 7:05 PM with an initial luminosity of 159E30.

At 7:35 PM, Operations tried to stash antiprotons, but the first transfer was lost; it never made it into the AP3 transfer line. Operations reported at 8:28 PM that the extraction kicker shutter was closed (A:ESHTST). It had closed for the reverse proton tune up prior to stashing and never reopened as it should have.

The Recycler successfully completed the last of six stashes for the day at 9:06 PM.

Linac experts completed their repairs on KRF3 at 9:21 PM.

Sunday November 12

The midnight shift began with Operations monitoring store 5063 and a stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 6:14 AM, Pbar suffered the first of many Debuncher RF alarms (DRF1-8) over the next twenty minutes with an ENI Not OK indication. Operators replaced the ENI at 7:03 AM.

At 10:21 AM, a MI RF station (MIRF14) tripped off. The station tripped off after each attempt to reset. Operators bypassed the station. This reduced the intensity MI was able deliver. Operations called in experts.

The RF expert reported at 11:23 AM that he found an LCW leak dripping on equipment for MIRF14. The Duty Mechanic reported at 12:15 PM that he had rerouted the LCW leak away from the equipment. The leak will be repaired later. Operations increased intensity to normal.

At 1:16 PM, a Booster RF station (BRF5) tripped off and would not reset. Operators bypassed the station.

At 4:17 PM, Operations reported that there was a LCW leak somewhere in Pbar. Operators searched all the service buildings; they found no signs of a leak, which meant the leak was probably in the tunnel.

Operations reported the following status at 8 PM: store 5063 luminosity = 29.83E30, Pbar stack = 34.8E10, Recycler stash = 295E10.

Operations terminated store 5063 at 9:03 PM.

Operations prepared D0 for an access at 9:22 PM.

Cryo system techs began work on the TeV sector A3 wet engine at 9:34 PM. They had the work completed 30 minutes later.

The Recycler completed the eighth and last stash for the day at 10:36 PM.

D0 came out of access at 10:50 PM.

Operations put the TeV through a dry squeeze at 11:31 PM.

Operations put the TeV into shot setup at 11:54 PM.

Monday November 13

The midnight shift began with the TeV in shot setup, Operations monitoring a stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations established store 5065 at 2:32 AM with an initial luminosity of 200.34E30.

The Recycler completed the third and last stash for the midnight shift at 6:55 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. Experts will access Pbar this afternoon to look for the LCW leak.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/13/2006    email Fermilab